[The following information applies to the questions displayed below.]

On January 1, 2021, Weaver Corporation purchased a patent for $267,000. The remaining legal life is 20 years, but the company estimates the patent will be useful for only six more years. In January 2023, the company incurred legal fees of $87,000 in successfully defending a patent infringement suit. The successful defense did not change the company’s estimate of useful life. Weaver Corporation’s year-end is December 31.

a. Record the purchase in 2021; amortization in 2021; amortization in 2022; legal fees in 2023; and amortization in 2023. (If no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field.)

1. January 1, 2021 - Record the patent purchase.

2. December 31, 2021 - Record amortization on the patent.

3. December 31, 2022 - Record amortization on the patent.

4. January 1, 2023 - Record legal fees on the patent infringement suit.

5. December 31, 2023 - Record amortization on the patent

b.  What is the balance in the Patent account at the end of 2023?
Answer #1
Date Account Tittle Debit Credit
2021 $ $
Jan 01 Patent 267,000 267,000
Cash
Patent Purchased
Dec. 31 Amortization Expense 44,500
Accumulated Amortization-Patent 44,500
Annual amortization of patent for useful life-6 Yr
2022
Dec. 31 Amortization Expense 44,500
Accumulated Amortization- Patent 44,500
Annual amortization of patent
2023
Jan. 01 Patent 87,000
Cash 87,000
Legal fees on patent infringement suit
Dec. 31 Amortization Expense 66,250
Accumulated Amortization 66,250
Annual amortization of patent
b) Balance of Patent at the end of 2023, Dec. 31
Patent Cost ( Purchase Price + Legal Fees) 354,000
Less: Amortization Expense-Patent (155,250)
198,750
Notes: Calculation of Amortization Expense.
Although legal life of patent is 20 year but useful life is only 6 years. The Patent cost
267,000 shall be amortized only in 6 years. So, Annual amortization cost will be
267,000/6, $44,500.
The Patent value after 2 years of amortization is 267,000 - 89,000 = 178,000. The
legal fees incurred in defending law suit is 87,000. Thus, the cost value due to this
is 178,000 + 87,000= 265,000 and the remaining life is now 4 year. So, Annual
amortization expense for 3rd year is 265,000/4= $66,250